# Dressed to Gamble: How Poker Drives the Dynamics of Wearables and Visits on Decentraland’s Social Virtual World

### v1

@@ Abstract
Decentraland is a blockchain-based social virtual world touted to be a creative space owned by its community, unlike previous virtual worlds. Its users can create and publish wearables, virtual garments to customize avatars, which can be then sold or given away via the blockchain. Decentral Games (DG), a single project owning two prominent in-world casinos, has by far created the most wearables, with these being necessary to earn cryptocurrency in their flagship game ICE Poker. We thus present a comprehensive study that investigates how DG and ICE Poker influence the overall dynamics of Decentraland wearables and in-world visits. To this end, we analyzed 5.9 million wearable transfers made on the Polygon blockchain (and related sales) over a two-year period, and 677 million log events of in-world user positions in an overlapping 10-month period. We found that the influence of DG and Ice Poker is not only significant, but also substantial for transfers and sales monetary value of wearables, and very large for daily unique visitors and time spent in the virtual world. Despite several alternative in-world economic and artistic initiatives in Decentraland, some of which have attracted much attention from the general public, online poker appears to be the main driver of the analyzed dynamics. Our work thus contributes to the current understanding of user behavior in social virtual worlds and it is among the first to study the emerging phenomenon of blockchain-based online gambling in virtual spaces.
@@ 1 Introduction
In recent years, there has been much discussion around the Metaverse, an envisioned fully immersive iteration of the Internet, composed of a shared collection of interoperable 3D virtual worlds. The Metaverse remains a vision that is in its early stages Cheng et al. (2022), but many social virtual worlds usually associated with it thrived during the Covid-19 pandemic Kerdvibulvech (2022). In particular, interest surged in the wake of Facebook’s rebranding into Meta in late 2021, which signaled the company’s quest to realize such an ambitious vision Kraus et al. (2022). In these virtual worlds, interaction with the environment and with other people is done primarily through an avatar, an embodied representation of the user. Further, avatar customization (e.g., via virtual wearables) is a significant in-world factor of purchase intent, as well as one of the main pleasures for users Bleize & Antheunis (2019). However, it has been argued Zhou et al. (2018) that many of the ownership policies for in-world digital goods —such as those used in avatar customization— are detrimental to users due to the imbalance in terms of rights between platforms and users. For example, traditionally digital assets are not transferable independently of the platform; hence, these assets remain locked-in inside the virtual world or are forever lost when the platform ceases to function.
In response, blockchain technology has been heralded by many tech enthusiasts as a decentralized and transparent approach to digital property in virtual worlds Cannavo & Lamberti (2020). Arguably, Decentraland is the earliest and most popular blockchain-based virtual world, touted to be the first one “built, governed, and owned by its users” in the project’s official website. In Decentraland, users can transfer in-world land, usernames, and avatar wearables, either as part of a sale or for free. All of these transfers are recorded on the blockchain Ethereum, or in the case of wearables on Polygon, a more scalable and compatible blockchain. Hence, there is a trove of public data regarding the dynamics of this social virtual world.
@@
Therefore, herein we delve into how DG and ICE Poker influence the dynamics of ownership transfers of wearables and in-world visits in Decentraland, a virtual world that touts its social and creative aspects of all kinds, but that seems to be driven by gambling. In particular, we seek to answer the following research questions (RQs):
@@ 2 Background
In this section we briefly describe the main mechanisms, jargon, and events of both Decentraland and DG referred to in the related work, analyses, and discussion thereof.
@@ 2.1 Decentraland
Publicly released in February 2020, Decentraland’s social virtual world revolves around the so-called tokenomics, i.e., a decentralized economy based on the exchange of assets via cryptographic tokens on the blockchain Lo & Medda (2020). On Ethereum and Polygon, these exchanges are made from one address to another address via smart contracts, which set the automatic rules of the transactions. These addresses identify either an account or a smart contract that can participate on the blockchain.
Decentraland has its own cryptocurrency, called MANA, with which users can buy in-world digital assets, such as parcels of land and wearables. The proof of ownership of these assets is in the form of non-fungible tokens (NFTs) —unique and non-interchangeable, unlike the fungible cryptocurrencies. Possession of MANA is also used as the basis of voting power for platform decisions via the decentralized autonomous organization (DAO) of Decentraland, a form of blockchain-based governance without a central authority Goldberg & Schär (2023).
@@ 2.2 Decentral Games
In October 2021, DG released ICE Poker, touted to be a Metaverse play-to-earn game, i.e., one in which users are able to earn cryptocurrency simply by playing. In preparation, DG also launched the cryptocurrency ICE to function as the in-game currency for rewards and minting wearables. The following month, DG also published several exclusive ICE-themed wearables that granted attractive bonuses while playing the game, based on their ranking level. These were made available for minting in three batches to users that had a certain amount of DG and ICE.
@@ 3 Related Work
While we are still in the early stages of Metaverse development, numerous independent virtual worlds have already sprung up over the past few decades. Notable examples include Second Life, Meta Horizon Worlds, and massively multiplayer online (MMO) games such as World of Warcraft, Minecraft, Roblox and Fortnite. However, these virtual environments have faced significant criticism regarding the governance and ownership of in-world assets. In Second Life, for instance, there has been a perceived emergence of a new form of virtual feudalism Grimmelmann (2009). World of Warcraft has grappled with conflicts between the game’s publisher and players over agency and ownership of in-game assets Glas (2010). Additionally, there has been a general separation between content and platform ownership in virtual worlds, often to the detriment of users, whose rights over bought or created in-platform content are usually restricted Zhou et al. (2018).
@@
Recently, the fashion industry’s attention has shifted towards popular MMO games, particularly through the creation of branded skins—cosmetic options for characters, with an increasing interest in newer blockchain-based virtual worlds Rodriguez Sanchez & Garcia-Badell (2022). Notably, several prestigious luxury brands are venturing into NFT collectibles Joy et al. (2022), participating in events like the Metaverse Fashion Week or designing wearables for Decentraland Gonzalez (2022). Among the scarce scholarly investigation into Decentraland’s wearables, we find a quantitative analysis in which it emerged that they exhibit peculiar dynamics Trujillo & Bacciu (2023). For instance, wearables are mainly distributed for free to boost engagement in other cryptoasset or Metaverse initiatives. Only a minor portion of them are actually sold on the platform’s marketplace. Furthermore, the price of these wearables is largely influenced by their pre-established rarity, which escalates the cost even further for unique items (i.e., having a set minting limit of one). The study also identified a specific account that stands out in terms of wearable items collections, having published more than 80 at the time of the study. This is noteworthy, given that over half of all wearable creators in Decentraland typically author only a single collection. This creator is associated with Decentral Games, a company whose primary offering is a poker gaming experience within Decentraland virtual casinos.
Online gambling is a topic of research in its own right. This includes studies on the legality of online gambling and cryptocurrencies in different jurisdictions Zborowska et al. (2012); Yeoh (2017), the role of virtual communities in either motivating gamblers or protecting them from excessive behaviour Sirola et al. (2021); Savolainen et al. (2022), the overlap of personality and demographic traits among cryptocurrency traders and problem gamblers Johnson et al. (2023); Delfabbro et al. (2021), how those traits are partly different in professional gamblers Newall & Talberg (2023), and how playing habits in low-stakes gambling might influence the perception of financial risk in general Armstrong et al. (2018).
@@ 4.2 Data collection
For data on DG wearable upgrades, we again used Alchemy to retrieve the log events for the period of study of DG’s smart contract which manages the respective on-chain transactions, called “iceRegistrant”. We then used the aforementioned wearable metadata to derive the ranking level and rarity limit of the upgraded tokens.
For in-world user tracking, we utilized the corresponding data archives from Atlas Data Corp, a blockchain analytics company that obtained a grant from Decentraland’s DAO to finance a system to keep record of user sessions. These data are log events (created every circa 20 seconds) of the current location of every connected user on the several servers that manage the 3D rendering of in-world scenes and avatars on Decentraland. Given that this system started operating in mid July 2022, we narrowed our analysis of in-world visits to only 10 months, from August 2022 to May 2023. Moreover, we take into consideration only users with an Ethereum-compatible address, that is, we discard the small minority of guest users not associated with an account. We then transformed the more than 677 million log events for this time frame into sessions with a start and end for each land parcel and account address. Finally, we segmented the respective sessions’ duration on a daily basis, adding 10 seconds to each to compensate for possible shortenings due to the 20-second polling rate of the log events.
@@ 5 Analyses and Results
For the two-year period of study, a total of 4,629 collections with 8,450 wearable items (i.e., designs) were published by 1,517 distinct creators. Of these, we identified and manually verified two accounts that created all of the DG wearables, based on marketplace offerings, item names, and user profiles on Decentraland. Both DG accounts (0.1% of creators) published 117 collections with 723 items (8.6% of the total available items). Henceforth, we refer to these as DG wearables.
@@ 6 Discussion
Based on our results, we can respond affirmatively to both RQ1 and RQ2: wearable transfers on Decentraland are indeed disproportionately related to those created by DG, and the ICE Poker casinos do receive significantly (and remarkably) more visits than other in-world locations. In the following, we first discuss our main findings regarding the influence of DG and ICE Poker on Decentraland, we comment on the in-world economic mechanisms that allowed such influence to exist, we reflect upon the ambiguous legal standing of ICE Poker, and, finally, we describe the main limitations of our study.

### v2

@@ Abstract
Decentraland is a blockchain-based social virtual world touted to be a creative space owned by its community. In it, users can publish wearables used to customize avatars, which can be then sold or given away via blockchain transfers. Decentral Games (DG), a single project owning prominent in-world casinos, has by far created the most wearables, necessary to earn cryptocurrency in its flagship game ICE Poker. Herein, we present a comprehensive study on how DG and ICE Poker influence the dynamics of wearables and in-world visits in Decentraland. To this end, we analyzed 5.9 million wearable transfers made on the Polygon blockchain (and related sales) over a two-year period, and 677 million log events of in-world user positions in an overlapping 10-month period. We found that the platform-wise number of transfers and sales monetary value of wearables were disproportionally related to DG, and that its two ICE Poker casinos (less than 0.1% of the world map) represented a very large average share of daily unique visitors (33%) and time spent in the virtual world (20%). Despite several alternative in-world economic and artistic initiatives in Decentraland, some of which have attracted much attention from the general public, a single third-party online poker game appears to be the main driver of the analyzed dynamics. Our work thus contributes to the current understanding of user behavior in social virtual worlds, and it is among the first to study the emerging phenomenon of blockchain-based online gambling in virtual reality spaces.
@@ 1 Introduction
In recent years, there has been much discussion around the Metaverse, an envisioned fully immersive iteration of the Internet, composed of a shared collection of interoperable 3D virtual worlds. The Metaverse remains a vision that is in its early stages Cheng et al. (2022), but many social virtual worlds usually associated with it thrived during the Covid-19 pandemic Kerdvibulvech (2022). In particular, interest surged in the wake of Facebook’s rebranding into Meta in late 2021, which signaled the company’s quest to realize such an ambitious vision Kraus et al. (2022). In these virtual worlds, interaction with the environment and with other people is done primarily through an avatar, an embodied representation of the user. Further, avatar customization (e.g., via virtual wearables) is a significant in-world factor of purchase intent, as well as one of the main pleasures for users Bleize & Antheunis (2019). However, it has been argued Zhou et al. (2018) that many of the ownership policies for in-world digital goods —such as those used in avatar customization— are disadvantageous to users. This is primarily due to the imbalance in rights between the users and the proprietors of the virtual world platforms. For example, traditionally digital assets are not transferable independently of the platform; hence, these assets remain locked-in inside the virtual world or are forever lost when the underlying platform ceases to function.
In response, blockchain technology has been heralded by many tech enthusiasts as a decentralized and transparent approach to digital property in virtual worlds Cannavo & Lamberti (2020). Arguably, Decentraland is the earliest and most popular blockchain-based virtual world, touted to be the first one “built, governed, and owned by its users” in the project’s official website. In Decentraland, users can transfer in-world land, usernames, and avatar wearables, either as part of a sale or for free. All of these transfers are recorded on the blockchain Ethereum, or in the case of wearables on Polygon, a compatible and more scalable blockchain. Hence, there is a trove of public data regarding the dynamics of this social virtual world.
@@
Therefore, intrigued by both the peculiar behavior of the aforementioned account managed by DG and the economic incentive to play in its casinos, herein we delve into how DG and ICE Poker influence the dynamics of ownership transfers of wearables and in-world visits in Decentraland. In particular, we seek to answer the following research questions (RQs):
@@ 2 Background
In this section we briefly describe the main mechanisms, jargon, and events of both Decentraland and Decentral Games referred to in the related work, analyses, and discussion thereof.
2.1 Blockchain technologies
Blockchains are decentralized, immutable ledgers that facilitate secure and transparent recording of transactions across a network of computers. They are the technological backbone of cryptocurrencies such as Bitcoin (BTCNakamoto (2008) and Ether (ETHWoodet al. (2014), digital currencies that enable financial transactions without the need for a central authority like a bank. Blockchains also provide the underlying mechanisms for non-fungible tokens (NFTs), unique identifiers often used to represent ownership of digital objects (e.g, art and collectibles) Nadini et al. (2021). Unlike fungible cryptocurrencies, NFTs are unique and non-interchangeable. Furthermore, blockchains enable decentralized autonomous organizations (DAOs), a form of token-based governance without a central authority, allowing for distributed decision-making and management Hassan & De Filippi (2021).
The Ethereum blockchain Woodet al. (2014), in particular, is a versatile platform that supports all of the aforementioned applications, utilizing smart contracts —self-executing programs that automatically enforce the terms of an agreement. Decentraland’s features are based on the Ethereum blockchain, except for transactions concerning wearables; these are instead based on Polygon Kanani et al. (2021), a higher layer scaling solution with its own cryptocurrency (MATIC); it is compatible with Ethereum and is designed to improve transaction speeds and reduce costs while maintaining security and decentralization.
@@ 2.2 Decentraland
Publicly released in February 2020, Decentraland’s social virtual world revolves around the so-called tokenomics, i.e., a decentralized economy based on the exchange of assets via cryptographic tokens on the blockchain Lo & Medda (2020). On Ethereum and Polygon, these exchanges are made from one address to another address via the aforementioned smart contracts, which set the automatic rules of the transactions. These addresses identify either an account or a smart contract that can participate on the blockchain.
Decentraland has its own cryptocurrency, called MANA, with which users can buy in-world digital assets, such as parcels of land and wearables. The proof of ownership of these assets is in the form of NFTs. Possession of MANA is also used as the basis of voting power for platform decisions via the decentralized autonomous organization of Decentraland Goldberg & Schär (2023).
@@ 2.3 Decentral Games
In October 2021, DG released ICE Poker, touted to be a Metaverse Play-to-Earn (P2E) game, i.e., one in which users are able to earn cryptocurrency simply by playing. In preparation, DG also launched the cryptocurrency ICE to function as the in-game currency for rewards and minting wearables. The following month, DG also published several exclusive ICE-themed wearables that granted attractive bonuses while playing the game, based on their ranking level. These were made available for minting in three batches to users that had a certain amount of DG and ICE.
@@ 3 Related Work
While we are still in the early stages of Metaverse development, numerous independent virtual worlds have already sprung up over the past few decades. Notable examples include Second Life, Meta Horizon Worlds, and massively multiplayer online (MMO) games such as World of Warcraft, Minecraft, Roblox and Fortnite. However, these virtual environments have faced significant criticism regarding the governance and ownership of in-world assets. In Second Life, for instance, there has been a perceived emergence of a new form of virtual feudalism Grimmelmann (2009), while Roblox has been criticised for profiting from child labor Parkin (2022) and for allowing harmful content and games with gambling-like mechanisms in its user-generated virtual worlds Kou & Gui (2023). World of Warcraft has grappled with conflicts between the game’s publisher and players over agency and ownership of in-game assets Glas (2010). Additionally, there has been a general separation between content and platform ownership in virtual worlds, often to the detriment of users, whose rights over bought or created in-platform content are usually restricted Zhou et al. (2018).
@@ 3.1 Fashion in Virtual Worlds
Recently, the fashion industry’s attention has shifted towards popular MMO games, particularly through the creation of branded skins—cosmetic options for characters, with an increasing interest in newer blockchain-based virtual worlds Rodriguez Sanchez & Garcia-Badell (2022). Notably, several prestigious luxury brands are venturing into NFT collectibles Joy et al. (2022), participating in events such as the Metaverse Fashion Week or designing wearables for Decentraland Gonzalez (2022), as well as retailing and engaging in other kinds of customer engagement within Decentraland’s social virtual world Goldberg et al. (2023).
In fact, Decentraland’s wearables are mainly distributed for free to boost engagement in other cryptoasset or Metaverse initiatives, with only a minor portion of them (3.4%) being sold on the platform’s marketplace Trujillo & Bacciu (2023). Besides wearables, other kinds of NFTs available within the virtual world (e.g., tickets, memberships) are likewise transferred for free as part of promotional activities instead of being part of a sale transaction Goldberg et al. (2023). Perhaps unsurprisingly, when sold in Decentraland’s marketplace, the price of wearables is largely influenced by their pre-established rarity Trujillo & Bacciu (2023), which escalates even further for unique items —i.e., having a set minting limit of one. Incidentally, over half of all wearable creators in Decentraland typically author only a single collection (mostly with a single item), but as mentioned in the Introduction, a specific account related to Decentral Games stands out having published circa 80 collections with more than 600 items total in a year and a half of activity Trujillo & Bacciu (2023). These wearables are the only way to get access to the games played on the company’s casinos within Decentraland, with the primary offering being ICE Poker, a virtual poker gaming experience.
3.2 Poker and Online Gambling
The chance-based nature of the game of poker appeals to a wide audience of casual players, while the element of skill, particularly evident over many hands, entices professional players Fiedler & Rock (2009). The introduction of Internet to gambling activities has provided easy access for the general public, round-the-clock access, shorter intervals between bets, and instant reinforcements Griffiths & Barnes (2008). After 2003, online poker experienced tremendous growth, with a total market value reaching 3 to 4 billion USD by 2010, with approximately 5.5 million players participating in virtual poker rooms across the globe Fiedler & Wilcke (2011). Among these players, a subset has emerged as professionals, especially adept at multitabling —simultaneously playing multiple games to further reduce the element of chance. As reliance on poker earnings grew, these players developed a strong extrinsic motivation to play, even if not entirely removed from the gameplay. Further, it has been found that playing for real money strongly influences both game dynamics and player experiences in poker games Zaman et al. (2014).
Notably, professional gamblers share psychological traits with problem gamblers, highlighting the complex interplay between skill and risk Newall & Talberg (2023). Problem gamblers, often drawn to low-stakes gambling, may develop distorted risk perceptions and ultimately underestimate financial risks in general Armstrong et al. (2018). Interestingly, players are socially motivated to gamble by the virtual communities they form, while these same communities also serve as safeguards against excessive behavior Sirola et al. (2021); Savolainen et al. (2022). Online poker’s unique subculture —which includes websites, forums, and chat rooms from all around the world— fosters cooperative behavior among players through the sharing of information and strategies. Nonetheless, the game’s inherent competitiveness persists and is reflected in the hierarchical structures within these global communities O’Leary & Carroll (2013).
Despite online gambling being a global phenomenon, its legality varies across jurisdictions, thereby posing significant challenges for both players and operators as they attempt to determine whether a specific game is legal to play or not Zborowska et al. (2012). The advent of blockchain technology has ignited enthusiasm due to its potential to improve fairness and transparency in the sector, promising decentralized and autonomous solutions for regulators and operators alike. However, at the same time it ultimately adds complexity to the legal framework of online gambling, due in great part to the current regulatory uncertainties of cryptocurrencies and other blockchain-based assets Gainsbury & Blaszczynski (2017).
3.3 Crypto Asset Trading
Cryptocurrency trading has emerged as a disruptive force in the financial landscape since the inception of Bitcoin in 2008. While market capitalizations have reached trillions of dollars, the trading environment remains characterized by high volatility, often leading to substantial crashes followed by significant rebounds Fang et al. (2022); OECD (2022). NFTs entered the scene in 2014, gaining widespread attention in early 2021 due to very high-profile sales. Digital art, music albums, and even tweets and sports highlights have been registered as unique crypto assets and subsequently sold for millions of dollars each, causing considerable hype Houser & Holden (2022). NFT users exhibit a profound enthusiasm for the foundational technologies that facilitate unique avenues for creative expression and the development of innovative content creation business models. However, these positive sentiments are counterbalanced by the contentious matter of digital content ownership, the prevalence of low-quality NFTs, the risk of scams, potential money laundering activities, and the evolving framework of legal measures Sharma et al. (2022).
The literature also shows an overlap of personality and demographic traits among problem gamblers and cryptocurrency traders Johnson et al. (2023); Delfabbro et al. (2021b). In both cryptocurrency trading and gambling, decisions are often based on limited information, short-term motives for gain, and with high uncertainty; hence, people attracted to latter are likely to be attracted to the former Delfabbro et al. (2021a). Numerous operators in the crypto gambling sector provide access to a variety of betting options, games, and online casinos. However, these platforms often pose a significant risk as they are easily accessible to minors and vulnerable individuals, and typically offer minimal to no consumer protection features. The legal landscape for these platforms is also a concern, as they remain largely unregulated in many countries Andrade et al. (2023).
Gambling aside, the legality of trading both cryptocurrencies and NFTs also varies greatly across jurisdictions. Some countries embrace these digital assets, while others impose strict regulations. Factors include tax treatment, anti-money laundering compliance, and investor protection. Researchers and policymakers continue to grapple with these complexities, seeking a balance between industry innovation and safeguarding the public Yeoh (2017).
3.4 Play-to-Earn Games
The emergence of blockchain technologies has also given rise to a new form of monetized online gaming known as “Play-to-Earn” (P2E) games. In P2E games, players receive monetary rewards for their gameplay, mainly in the form of the game’s specific cryptocurrency. Axie Infinity —a Pokémon-style battle game with collectible axolotl-like creatures— is arguably one of the most well-known examples, making use of cryptocurrencies and NFTs for its in-game economy. Its popularity peaked in 2021, particularly in the Philippines and later Venezuela, where it represented a viable source of earnings for low-income people during the height of the pandemic De Jesus et al. (2022).
However, extrinsically motivated gameplay, i.e., mainly driven by external rewards such as financial gain, can reduce intrinsic motivation and consequently have negative implications for player experience Delic & Delfabbro (2024); Hong et al. (2023). A survey conducted with a group of Axie Infinity players in 2020–21 has shown that, while some players enjoyed social interactions, overall intrinsic motivation tended to be low, causing high amounts of stress and unwanted tension to reach daily quotas De Jesus et al. (2022). This assessment was later corroborated by a study that analyzed online conversations about the game Delic & Delfabbro (2024).
Interestingly, part of the initial success of Axie Infinity was due to delegation models that emerged organically within the community, called scholarships De Jesus et al. (2022). These models allow players to share their knowledge or delegate their gameplay (and assets) to others, in exchange for a share of the rewards. Nonetheless, scholarship models and related delegation mechanisms within P2E games are potentially exploitative. While P2E games offer players from less-advantaged countries the prospect of earning income above the national average wage, players might find themselves participating in a scheme that provides low-cost labor for wealthier actors who hold valuable NFTs Delic & Delfabbro (2024). Curiously, ICE Poker integrates by design some elements of P2E games, offering the possibility to earn tokens just by participating in certain events, and a delegation mechanism similar to the community-based scholarships in Axie Infinity.
@@ 4.2 Data collection
For data on DG wearable upgrades, we again used Alchemy to retrieve the log events for the period of study of DG’s smart contract on Polygon [^2] Address 0xC9a67eD1472A76d064C826B54c144Ca00DAE4015 which manages the respective on-chain transactions, called “iceRegistrant”. We then used the aforementioned wearable metadata to derive the ranking level and rarity limit of the upgraded tokens.
For in-world user tracking, we utilized the corresponding data archives from Atlas Data Corp, a blockchain analytics company that obtained a grant  [^3] https://decentraland.org/governance/project/?id=7422a99e-8a25-4625-9b30-ab688de5dade from Decentraland’s DAO to finance a system to keep record of user sessions. These data are log events (created every circa 20 seconds) of the current location of every connected user on the several servers that manage the 3D rendering of in-world scenes and avatars on Decentraland. Given that this system started operating in mid July 2022, we narrowed our analysis of in-world visits to only 10 months, from August 2022 to May 2023. Moreover, we take into consideration only users with an Ethereum-compatible address, that is, we discard the small minority of guest users not associated with an account. We then transformed the more than 677 million log events for this time frame into sessions with a start and end for each land parcel and account address. Finally, we segmented the respective sessions’ duration on a daily basis, adding 10 seconds to each to compensate for possible shortenings due to the 20-second polling rate of the log events.
@@ 5.1 RQ1: Overall wearable transfers
From June 2021 to May 2023, a total of 4,629 collections with 8,450 wearable items (i.e., designs) were published by 1,517 distinct creators. Of these, we identified and manually verified two accounts [^4] Addresses 0x17a253c2ac0d5ba92cadbbf665e3390c9913dc5d and 0x00e5d44f6a296c10f159486f842838bd68f13e32 that created all of the DG wearables, based on marketplace offerings, item names, and user profiles on Decentraland. Both DG accounts (0.1% of creators) published 117 collections with 723 items (8.6% of the total available items). Henceforth, we refer to these as DG wearables.
@@ 6 Discussion
Based on our results, we can respond affirmatively to both RQ1 and RQ2: wearable transfers on Decentraland are indeed disproportionately related to those created by DG, and the ICE Poker casinos do receive significantly (and remarkably) more visits than other in-world locations. In the following, we first discuss our main findings concerning the influence of DG and ICE Poker on Decentraland, we comment on the initiatives and events affecting such influence over time, we reflect upon the ambiguous legal standing of ICE Poker, we explore the social implications of our findings, and, finally, we describe the main limitations of our study.